OpenStudy (anonymous):

How many solutions does this system of equations have? 5x - 4y = 11 4y = 5x - 9

OpenStudy (anonymous):

How many solutions does this system of equations have? 5x - 4y = 11=5x=4y+11 5x-4y-11=0 y=5x/4-11/4 x=4n+3 y=5n+1 y=1/4(5x-11)

OpenStudy (anonymous):

infinite. Correct?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Find two numbers whose sum is 42 and whose difference is 16

OpenStudy (anonymous):

can u help me in that one